hartnn (hartnn):

If cos(x+iy) cos(u+iv)=1,where x,y,u,v are real then show that tanh^2y cosh^2v=sin^2u

hartnn (hartnn):

expanding and comparing real and imaginary parts led to nowhere...

hartnn (hartnn):

@SithsAndGiggles

myininaya (myininaya):

@hartnn That is sin^2(u) not sinh^2(u) ?

hartnn (hartnn):

its sin^2 u

myininaya (myininaya):

The reason I'm asking is because ... \[\tanh^2(x) \cdot \cosh^2(x)=(\frac{e^x-e^{-x}}{e^x+e^{-x}})^2(\frac{e^x+e^{-x}}{2})^2=(\frac{e^x-e^{-x}}{2})^2=\sinh^2(x)\]

myininaya (myininaya):

I do realize we do have those different variables though

hartnn (hartnn):

typing mistakes in the question document is quite common, so if we can prove it to be sinh^2 u .....

myininaya (myininaya):

Well I'm not really sure yet about that. I'm just thinking.

hartnn (hartnn):

but i see no reason why it must not be sin^2 u

hartnn (hartnn):

@satellite73 @surjithayer @mathmale any ideas how to start...?

hartnn (hartnn):

@experimentX

OpenStudy (experimentx):

i think with usual cos(A+B) formula just note that cos(ix) = cosh(x) and sin(ix)=-sinh(x)

hartnn (hartnn):

Comparing real and imaginary parts cos x cosh y cos u cosh v-sin x sinh y sin u sinh v=1 cos x cosh y sin u sinh v+sin x sinh y cos u cosh v=0

OpenStudy (experimentx):

yes that is one of the tool needed to solve it.

hartnn (hartnn):

ok...then how would tanh terms and squares of them woud come into picture ?

OpenStudy (experimentx):

lol ... i haven't done it till now ... trolling on fb lemme try this and post in 10 min

hartnn (hartnn):

sure, thanks

OpenStudy (experimentx):

\[ \cos(x+iy) = \cos(x)\cosh(y) +i \sin(x)\sinh(y) \\ \cos(u+iv) = \cos(u)\cosh(v) + i \sin(u)\sinh(v) \] multiplying them gives, \[ \cos(x)\cosh(y) \cos(u)\cosh(v) - \sin(x)\sinh(y) \sin(u)\sinh(v) \\ + i (\cos(x)\cosh(y)\sin(u)\sinh(v)+\cos(u)\cosh(v)\sin(x)\sinh(y)) = 1 \]

hartnn (hartnn):

thats what i got

OpenStudy (experimentx):

the second part gives, \[ \tan(u)\tanh(v) = - \tan(x) \tanh(y) \] the first part gives, \[ \cos(x)\cosh(y) \cos(u)\cosh(v) - \sin(x)\sinh(y) \sin(u)\sinh(v) = 1\]

hartnn (hartnn):

right

OpenStudy (experimentx):

isn't there x term in question, looks like we migth begin by eliminating x term.

hartnn (hartnn):

no x term in the question, so can we try dividing the 2nd equation by cos x and replacing for tan x but there would still be sc x on right side...

OpenStudy (experimentx):

yes ... that's creating problem.

hartnn (hartnn):

any other approach..... like cos x+iy = sec u+iv or something...

OpenStudy (experimentx):

looks worth try

OpenStudy (experimentx):

but ... sec(a+ib) isn't going to be that nice

hartnn (hartnn):

even tan^2 a+ib ? cos^2 (x+iy) = 1+ tan^2 (u+iv) how about this ?

hartnn (hartnn):

still eliminating x would be a challenge in that expansion...

OpenStudy (experimentx):

IDK ... but where did you get the question BTW? did you check for the values of x,y,u,v where this relation is true? i couldn't guess any value other than 0's

hartnn (hartnn):

university question paper last year

hartnn (hartnn):

which means it isn't supposed to be hard :P

OpenStudy (experimentx):

haven't you already finished Masters?

hartnn (hartnn):

yep, i did solving previous year's question papers to help a friend

OpenStudy (experimentx):

damn man ... this topic is supposed to be covered undergrad. finding it hard to solve lol.

hartnn (hartnn):

exactly! there could be some twist...

OpenStudy (experimentx):

lemme me try this tonight. I am pretty distracted right now. If I solve it, i'll msg you on fb.

hartnn (hartnn):

sure, np., thanks :)

OpenStudy (experimentx):

i have also asked my friend to do it ... let's see what he gets.

myininaya (myininaya):

@hartnn I thought about this problem a lot yesterday and it stumps me too. I'm still trying but I'm not coming to anything useful.

hartnn (hartnn):

Even i tried solving it in different ways...no luck in google search too...this problem is bugging me from a week :O

myininaya (myininaya):

\[\text{ Suppose } \cos(x+iy) \cdot \cos(u+iv)=1 \text{ where x,y,u,v are real .} \\ \text{ W.T.S. } \tanh^2(y) \cdot \cosh^2(v)=\sin^2(u)\] I. \[\cos(x+iy)=\cos(x)\cosh(y)-i \sin(x) \sinh(y) \] II. \[\cos(u+iv)=\cos(u)\cosh(v)-i \sin(u) \sinh(v)\] Multiplying I and II gives: \[\cos(x)\cos(u)\cosh(y)\cosh(v)\\ -\sin(x)\sin(u)\sinh(y)\sinh(v) \\ -i \cos(x)\sin(u)\cosh(y)\sinh(v) \\ -i \sin(x) \cos(u) \sinh(y) \cosh(v) \] Now I times II is suppose to give 1 since we are supposing this is true for real values x,y,u, and v. So we have the equation: \[\cos(x)\cos(u)\cosh(y)\cosh(v)\\ -\sin(x)\sin(u)\sinh(y)\sinh(v) \\ -i \cos(x)\sin(u)\cosh(y)\sinh(v) \\ -i \sin(x) \cos(u) \sinh(y) \cosh(v) =1\] This means by comparing real and imaginary parts we have two equations: \[\ \text{ A. } cos(x)\cos(u)\cosh(y)\cosh(v)-\sin(x)\sin(u)\sinh(y)\sinh(v)=1 \\ \text{ and } \\ \text{ B. }\cos(x)\sin(u)\cosh(y)\sinh(v)+\sin(x)\cos(u)\sinh(y)\cosh(v)=0\] So let's remind ourselves of what we want to show: \[\tanh^2(y) \cosh^2(v)=\sin^2(u)\] So we are suppose to take equations A. and B. and somehow eliminate the x part and basically try to get sin^2(u) on one side of an equation. Since we have the square thing on both sides, I wonder if that means we need to multiply A and B... Let's try... A*B=1*0=0 So we have: \[\cos^2(x)\sin(u)\cos(u)\cosh^2(y)\cosh(v)\sinh(v)\\ +\sin(x)\cos(x)\cos^2(u)\sinh(y)\cosh(y)\cosh^2(v) \\ - \sin(x)\cos(x)\sin^2(u)\sinh(y)\cosh(y)\sinh^2(v) \\ -\sin^2(x)\sin(u)\cos(u)\sinh^2(y)\sinh(v)\cosh(v)=0\] I think we may be able to do some factoring here... This is what happens if we group 1 and 4 together then 2 and 3. (talking about the terms) \[\sin(u)\cos(u)\sinh(v)\cosh(v) \cdot [\cos^2(x)\cosh^2(y)-\sin^2(x)\sinh^2(y)] \\ +\sin(x)\cos(x)\sinh(y)\cosh(y) \cdot [\cos^2(u)\cosh^2(v)-\sin^2(u)\sinh^2(v)]=0\] \[\sin(u)\cos(u)\sinh(v)\cosh(v)[\cos^2(x)\cos^2(y)-\sin^2(x)\sinh^2(y)]=\\-\sin(x)\cos(x)\sinh(y)\cosh(y)[\cos^2(u)\cosh^2(v)-\sin^2(u)\sinh^2(v)]\] I'm going to go ahead and post this just in case my openstudy dies on me. It is probably not helpful because I have been staring at this last thing I wrote forever.

myininaya (myininaya):

I thought I was going to be able to use cosh^2(x)-sinh^2(x)=1 :( but then those lousy factos cos^2(x) and sin^2(x) got in the way

hartnn (hartnn):

thanks for the link @ganeshie8 i would post that link so that experimentX and myin would get the solution too https://docs.google.com/file/d/0B7ss_Y8M_VEgMzRYVmdidjFnN00/edit it was truly difficult problem....unexpected to come in university exam :O

myininaya (myininaya):

Yeah, I don't think I would have came up with that solution.